# On residual finiteness of graphs of free groups with cyclic edge groups

### v1

@@ 4. Detecting very unbalanced elements
Let $A\subset F$ be minimal so that for each $g\in F$ represented by a $u_{j}$ or $v_{j}$, either $g$ or $g^{-1}$ is conjugate to a (unique) element of $A$. Clearly, $|A|\leq 2|J|$ and since the group elements represented by the $u_{j}$ and $v_{j}$ are not proper powers in $F$, $A$ is independent (Lemma 3.4). For $g\in A$, let $w_{g}$ denote a cyclically reduced word in the $\{a_{i}^{\pm}\}$ representing $g$. By the previous observation, the presentation can be modified so that every $u_{j}$ and every $v_{j}$ is equal to some $w_{g}$, $g\in A$. From now on, we will assume standard presentations are of this form. In particular, each standard presentation will come with a set $A$ as above.
@@ 5. Residual finiteness
Let $p$ be a path in $X(n,m)$ labeled by a word $w$ in $a$, $t$ representing a non-trivial element of $BS(1,q)$. Let $\hat{p}$ be a lift of $p$ to $\widehat{X}(n,m)$. If the sum of powers of $t$ in $w$ is non-zero, then for $m$ sufficiently large, $\hat{p}$ is a non-closed path in $\widehat{X}(n,m)$. If the sum of powers of $t$ in $w$ is zero, then $w$ represents a conjugate of a power of $a$. Indeed, using the relation $ta=a^{q}t$ one can transform $w$ into a word of the form $t^{-k}a^{l}t^{k}$ for some integers $k,l$ with $k\geq 0$. Thus for $n$ sufficiently large, $\hat{p}$ is not closed in $\widehat{X}(n,m)$. Since $w$ represented an arbitrary non-trivial element of $BS(1,q)$, this proves residual finiteness of $BS(1,q)$.
@@ 5.2. Building the cover
A group splitting as a finite graph of free groups with cyclic edge groups is residually finite if and only if it contains no very unbalanced elements. In particular, its residual finiteness can be checked algorithmically given a presentation of the splitting.
@@
Assume $g\in G$ is very unbalanced for $h,m,n$. As in the proof of Proposition 3.5, $g$ acts elliptically and $h$ acts loxodromically on the Bass-Serre tree of the HNN-extension. Let $v$ be a vertex fixed by $g$. Then $g^{n}=hg^{m}h^{-1}$ fixes the geodesic joining $v$ and $hv\neq v$. Since conjugates of edge groups are pairwise disjoint except when they are both the same conjugate of $\left\langle a\right\rangle$, the vertices $v$ and $hv$ are joined by a path of length at most $2$ stabilized by $g$ and there exists $h^{\prime}\in G$ and $p,q$ non-negative distinct integers satisfying the following: $h^{\prime}g^{m}h^{\prime-1}$ is a power of $ab^{p}$ and $h^{\prime}g^{n}h^{\prime-1}$ is a power of $ab^{q}$. Thus $ab^{p}$ and $ab^{q}$ share a common power, a contradiction since $p\neq q$ and $\left\langle a,b\right\rangle$ embeds in $G$. Therefore, $G$ has no very unbalanced elements.
@@
Is there a combinatorial property of the graph $\Gamma$ encoding when $G$ is solvable? linear? residually $p$-finite for a prime $p$?
More explicitly, each embedded cycle $C\to\Gamma$ determines a homomorphism $BS(m,n)\to G$ from a Baumslag-Solitar group to $G$, where $m,n$ are the loop-products of the cycle. We can associate to $G$ the collection $BS_{G}$ of Baumslag-Solitar groups appearing in this way. Some properties of $G$ are closely related to the properties of groups in $BS_{G}$. For example, $G$ is subgroup separable if and only if all groups of $BS_{G}$ are subgroup separable [Wis00]. And the work in this article implies that $G$ is residually finite if and only if all groups of $BS_{G}$ are residually finite. Thus one can expect that $G$ is residually $p$-finite (resp. solvable, linear) if and only if all groups of $BS_{G}$ are residually $p$-finite (resp. solvable, linear). Note that a Baumslag-Solitar group is linear if and only if it is residually finite.
We also expect an argument very similar to ours to characterize residual finiteness in graphs of special groups with cyclic edge groups, where a special group is the fundamental group of a compact special cube complex.

### v2

@@ 4. Detecting very unbalanced elements
Let $A\subset F$ be minimal so that for each $g\in F$ represented by a $u_{j}$ or $v_{j}$, either $g$ or $g^{-1}$ is conjugate to a (unique) element of $A$. Clearly, $|A|\leq 2|J|$ and since the group elements represented by the $u_{j}$ and $v_{j}$ are not proper powers in $F$, $A$ is independent. For $g\in A$, let $w_{g}$ denote a cyclically reduced word in the $\{a_{i}^{\pm}\}$ representing $g$. By the previous observation, the presentation can be modified so that every $u_{j}$ and every $v_{j}$ is equal to some $w_{g}$, $g\in A$. From now on, we will assume standard presentations are of this form. In particular, each standard presentation will come with a set $A$ as above.
@@
Remark 4.6.
In fact, each component of $\Gamma$ is the defining graph of a Generalized Baumslag-Solitar group, which embeds in $G$ as the commensurator of any vertex of the component. Every non-cyclic commensurator of an element of $G$ appears in this way. Clean components correspond exactly to the cases where the Generalized Baumslag-Solitar subgroup is unimodular in Levitt’s terminology ([Lev15]), i.e. LERF.
@@ 5. Residual finiteness
Let $p$ be a path in $X(n,m)$ labeled by a word $w$ in $a$, $t$ representing a non-trivial element of $BS(1,q)$. Let $\hat{p}$ be a lift of $p$ to $\widehat{X}(n,m)$. If the sum of powers of $t$ in $w$ is non-zero, then for $m$ sufficiently large, $\hat{p}$ is a non-closed path in $\widehat{X}(n,m)$. If the sum of powers of $t$ in $w$ is zero, then $w$ represents a conjugate of a power of $a$ (see the proof of Lemma 3.5). Thus for $n$ sufficiently large, $\hat{p}$ is not closed in $\widehat{X}(n,m)$. Since $w$ represented an arbitrary non-trivial element of $BS(1,q)$, this proves residual finiteness of $BS(1,q)$.
@@ 5.2. Building the cover
A group splitting as a finite graph of free groups with cyclic edge groups is residually finite if and only if all of its Baumslag-Solitar subgroups are residually finite. Morevoer, the residual finiteness can be determined algorithmically given a presentation of the splitting.
@@
Assume $g\in G$ is very unbalanced for $h,m,n$. By Lemma 3.4, $g$ acts elliptically and $h$ acts loxodromically on the Bass-Serre tree of the HNN-extension. Let $v$ be a vertex fixed by $g$. Then $g^{n}=hg^{m}h^{-1}$ fixes the geodesic joining $v$ and $hv\neq v$. Since conjugates of edge groups are pairwise disjoint except when they are both the same conjugate of $\left\langle a\right\rangle$, the vertices $v$ and $hv$ are joined by a path of length at most $2$ stabilized by $g$ and there exists $h^{\prime}\in G$ and $p,q$ non-negative distinct integers satisfying the following: $h^{\prime}g^{m}h^{\prime-1}$ is a power of $ab^{p}$ and $h^{\prime}g^{n}h^{\prime-1}$ is a power of $ab^{q}$. Thus $ab^{p}$ and $ab^{q}$ share a common power, a contradiction since $p\neq q$ and $\left\langle a,b\right\rangle$ embeds in $G$. Therefore, $G$ has no very unbalanced elements.
@@
Is there a combinatorial property of the graph $\Gamma$ encoding when $G$ is linear? residually $p$-finite for a prime $p$?
We also expect an argument very similar to ours to characterize residual finiteness in graphs of hyperbolic special groups with cyclic edge groups.